OpenStudy (anonymous):

find x and y, i is imaginary 3x-4iy=4+4i

OpenStudy (anonymous):

Assuming that x and y are real, we use the fact that a + bi = c + di if and only if a = c and b = d (where a, b, c, and d are real) So 3x = 4 -4y = 4